# Quantum Bayes’ rule and Petz transpose map from the minimal change principle

### v2

@@ Abstract
Bayes’ rule, which is routinely used to update beliefs based on new evidence, arises from a principle of minimal change. This principle states that updated beliefs must be consistent with new data, while deviating minimally from the prior belief. Here, we introduce a quantum analog of the minimal change principle and use it to derive a quantum Bayes’ rule by minimizing the change between two quantum input-output processes, not just their marginals. This is analogous to the classical case where Bayes’ rule is obtained by minimizing the change between the joint input-output distributions. When the change is quantified by the fidelity, the quantum minimal change principle has a unique solution, and the resulting quantum Bayes’ rule recovers the Petz transpose map.
@@
Introduction.—Usually demonstrated by simple counting arguments involving urns and balls, Bayes’ rule has actually been argued to play a much deeper role in probability theory and logic, as the only consistent system for updating one’s beliefs in light of new evidence [1, 2, 3, 4, 5, 6]. As an alternative to the above axiomatic approach, Bayes’ rule can also be derived from a variational argument: the updated belief should be consistent with the new observations while deviating as little as possible from the initial belief. This is known as the minimal change principle [7, 8, 9, 10], which formalizes the intuition that the new information should be incorporated into the agent’s knowledge in the “least committal” way, e.g. without introducing biases unwarranted by the data. Such fundamental insights can be seen as at least a motivation, if not an explanation, for the extraordinary effectiveness of Bayesian statistical inference in virtually all areas of knowledge.
@@
If one considers quantum theory as a noncommutative extension of probability theory, one would expect that there would also be a sound analog of Bayes’ rule. However, the status of a quantum Bayes’ rule is still much debated, with many alternatives, often inequivalent to each other, having been proposed in the past decades [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Among these proposals, the Petz transpose map [22, 23] stands out as the only quantum Bayes’ rule that satisfies a set of axioms analogous to the classical Bayes’ rule [21].
@@
Attempts have also been made to derive a quantum Bayes’ rule from optimizations involving the “posterior” state. For instance, Ref. [15] minimizes a loss function related to the quantum relative entropy, while Refs. [24, 18] optimize an upper bound on a distance measure between two estimators of the initial and final states. However, these methods, while involving optimization, focus primarily on the marginals of the process other than the entire process. As a result, they do not fully reflect the minimality of change over the entire process, which is instead the core argument from which Bayes’ rule and its generalizations, such as Jeffrey’s theory of probability kinematics, emerge.
@@
In this work, we take a decisive step towards solving this problem by presenting a new approach to the quantum Bayes’ rule, based on a natural quantum analog of the minimal change principle, which involves the entire process, not just its marginals (see Fig. 1 for a schematic representation). Specifically, when the minimal change is quantified by the quantum fidelity [25, 26], the resulting quantum Bayes’ rule can be derived analytically and corresponds to the Petz transpose map in many cases. This connection further strengthens the link between Bayes’ rule, the minimum change principle, and the Petz transpose map, thus justifying their broader applications in physics, not only in quantum information theory, and possibly beyond.
@@
Classical Bayes’ rule from the minimal change principle.—Consider two random variables $X$ and $Y$ taking values in two finite alphabets, $\mathcal{X}=\{x\}$ and $\mathcal{Y}=\{y\}$, respectively. Assume that the initial belief about $X$ is represented by a distribution $\gamma(x)$, while the correlation between $X$ and $Y$ is given by a conditional probability distribution $\varphi(y|x)$. We can think of the latter as the “forward process” from $X$ to $Y$. Now, suppose we observe a certain value $Y=y_{0}$ and want to update our belief about $X$ in light of this new information. Bayes’ rule provides the formula for the posterior probability as:
@@
But how should the above be generalized to the situation where the new information does not come in the form of a definite value $y_{0}$ for $Y$, but as “soft evidence”, i.e. a probability distribution $\tau(y)$? As Jeffrey [27] argues, this is in fact the rule rather than the exception, since it is very rare that we can ascribe probability one to only one event $Y=y_{0}$, ruling out all others with absolute certainty. Jeffrey’s probability kinematics [27] and Pearl’s virtual evidence method [28] agree [29] that the correct way to proceed is to promote Bayes’ posterior to a full-fledged “reverse process” $\hat{\varphi}_{\gamma}(x|y)$ that yields the probabilities of $X$ given $Y$, so that the updated belief given the new evidence $\tau(y)$ becomes
@@
While Jeffrey’s and Pearl’s derivations of Eq. 2 rely on certain natural assumptions about the logic model or underlying Bayesian network, an alternative way to obtain Eq. 2 is through the minimal change principle [8, 7, 9, 10]. To use this principle, we consider the joint distribution of $X$ and $Y$ in the forward process, denoted as $P_{\rm fwd}(x,y):=\varphi(y|x)\gamma(x)$, and try to update it based on a subsequent observation of $Y$. The update will be done not only on the marginal distribution on $X$, but on the whole joint distribution $P_{\rm fwd}(x,y)$. The minimal change principle then determines the updated belief about the joint distribution, denoted as $P_{\rm rev}(x,y)$.
Concretely, the minimal change principle can be described as the minimization of a divergence measure between the prior belief $P_{\rm fwd}$ and the updated belief $P_{\rm rev}$ under the constraint that $P_{\rm rev}$ is compatible with the new observation $\tau(y)$. This can be expressed as the following optimization problem:
@@
where $\mathbb{D}\left(P_{\rm fwd},P_{\rm rev}\right)$ is a statistical divergence measure between the prior and posterior joint distributions. Common choices include the Kullback–Leiber (KL) divergence [7, 10], other $f$-divergences including Pearson divergence and Hellinger distance [30], zero-one loss [31], or the mean-square error of an estimation [32, 33]. In all these cases, the optimal value is achieved by
@@
The above equation shows that the minimal change principle applied to the joint distributions of $X$ and $Y$ leads to the Bayes–Jeffrey rule in Eqs. (1) and (2).
@@
Stochastic processes in the quantum case correspond to completely positive trace-preserving (CPTP) linear maps, also known as quantum channels. Following the classical Bayes’ rule, we start with a forward quantum process $\mathcal{E}$, i.e. a channel from system $A$ to system $B$. The goal is to determine a backward quantum process $\mathcal{R}$, which is again a channel, but going from $B$ to $A$. Note that the adjoint map $\mathcal{E}^{\dagger}$, defined by trace duality as $\operatorname{Tr}\!\left[\mathcal{E}(X)\;Y\right]=:\operatorname{Tr}\!\left[X\;\mathcal{E}^{\dagger}(Y)\right]$ for all operators $X$ and $Y$, is a completely positive linear map from $B$ to $A$, although it is generally not trace-preserving and thus not a channel. As in the classical case, we assume a prior belief about the input, but instead of a probability distribution, the prior belief is now given as a quantum state $\gamma\in\mathcal{S}({\mathcal{H}}_{A})$.
A quantum analog of the joint input-output distribution of a stochastic process, such as the above $P_{\rm fwd}(x,y)=\varphi(y|x)\gamma(x)$, is what is known as a quantum state-over-time function [34, 20]. This function, which for channel $\mathcal{E}$ and input state $\rho$ is generically denoted as
$\mathcal{E}\star\rho\;,$
produces an operator on the joint Hilbert space ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$ that represents the process of $\mathcal{E}$ acting on $\rho$ as its initial state. Similar to a joint distribution, the marginal operators of $\mathcal{E}\star\rho$ on system $A$ and system $B$ corresponds to the input state $\rho$ and the output state $\mathcal{E}(\rho)$, respectively.
Here, we adopt a representation of quantum processes proposed in [35, 36], which is a variation of the Liefer–Spekkens state over time [16] up to a partial transpose, and is defined as
$Q_{\rm fwd}:=\mathcal{E}\star\gamma:=\left(\mathds{1}_{B}\otimes\sqrt{\gamma^{T}}\right)C_{\mathcal{E}}\left(\mathds{1}_{B}\otimes\sqrt{\gamma^{T}}\right),$ (6)
where $\mathds{1}_{B}$ is the identity on system $B$ and $C_{\mathcal{E}}=\sum_{i,j}\mathcal{E}(|i\rangle\!\!\!\;\langle j|)\otimes|i\rangle\!\!\!\;\langle j|\in\mathcal{L}_{+}({\mathcal{H}}_{B}\otimes{\mathcal{H}}_{A})$ is the Choi operator of the forward process $\mathcal{E}$ [37]. The superscript notation $\bullet^{T}$ denotes the transposition taken with respect to the orthonormal basis |$\ket{i}$ used to define the Choi operator. Computing the marginal operators we obtain
@@
We are thus using a representation that is a state-over-time function up to a partial transpose on system $A$. A benefit of this representation is that $Q_{\rm fwd}$ is always in $\mathcal{S}({\mathcal{H}}_{B}\otimes{\mathcal{H}}_{A})$, i.e., it is a bipartite state for all choices of channel $\mathcal{E}$ and prior state $\gamma$, and many divergence functions are well-defined and possess nice properties when evaluated on states.
@@
$Q_{\rm rev}:=(\mathcal{R}\star\tau)^{T}=(\sqrt{\tau}\otimes\mathds{1}_{A})\;C_{\mathcal{R}}^{T}\;(\sqrt{\tau}\otimes\mathds{1}_{A})\;.$ (8)
@@
The quantum minimal change principle.—We are now ready to formulate the quantum minimal change principle in a rigorous way. Here, we choose to measure “change” in terms of the (square-root) fidelity, defined as [25, 26]
@@
The quantum minimal change principle, similar to Eq. 5, minimizes the deviation between $Q_{\rm fwd}=\mathcal{E}\star\gamma$ and $Q_{\rm rev}=(\mathcal{R}\star\tau)^{T}$, or equivalently, maximizes their fidelity. The principle is formally expressed as follows:
$\begin{split}\max_{\mathcal{R}}\quad&F\left(\mathcal{E}\star\gamma,(\mathcal{R}\star\tau)^{T}\right)\;,\\ \text{subject to}\quad&\mathcal{R}\text{\rm~{}is CPTP}\;.\end{split}$ (11)
@@
Looking at this fact from another perspective, whenever $\mathcal{E}(\gamma)$ and $\tau$ commute, we find that our solution is consistent with the principles of Jeffrey’s probability kinematics, according to which the retrodictive channel that propagates the new information back to the input system should be independent of the new information. However, in the general quantum case where $\mathcal{E}(\gamma)$ and $\tau$ do not commute, our solution suggests an extension that deviates from the rules of Jeffrey’s probability kinematics. In any case, we find a strong connection between the quantum minimal change principle and Petz’s transpose map, which has already been argued to be a good candidate for a quantum Bayes’ rule [21].
@@
Before fleshing out the proof of the above theorem (the full details can be found in the Supplemental Material), let us first make some comments. First, it is easy to verify that the map given in Eq. 12 is indeed completely positive and trace-preserving. Complete positivity holds because $\mathcal{R}$ is the composition of three completely positive maps, that is, $D\bullet D^{\dagger}$, $\mathcal{E}^{\dagger}(\bullet)$, and $\sqrt{\gamma}\bullet\sqrt{\gamma}$. The trace-preservation condition is $\mathrm{Tr}[\mathcal{R}(\sigma)]=\mathrm{Tr}[D^{\dagger}\mathcal{E}(\gamma)D\sigma]=\mathrm{Tr}[\sigma]$ for any $\sigma$, equivalent to the condition
@@
    when $\mathcal{E}$ is a quantum-classical channel, e.g. a measurement channel, and the new information represented by $\tau$ is classical information about the measurement results.
The last condition in particular suggests a close relation with the theory of observational entropy and the corresponding second law [41, 38, 42].
Third, we see that the quantum analog of the Jeffrey-retrodicted distribution $\gamma^{\prime}(x)$ in Eq. 1 is the quantum state
@@
While the CPTP map $\mathcal{R}$ is always linear, the overall dependence of $\gamma^{\prime}$ on $\tau$ is not, in contrast to the classical case. Again, it is the commutativity condition $[\tau,\mathcal{E}(\gamma)]=0$ that restores the linear dependence of $\gamma^{\prime}$ on $\tau$. The functional dependence of the retrodicted state on the new information is an important aspect of the problem, which we leave open for future research, especially in view of the axiomatic approach proposed in [21], where linearity of the retrodiction is assumed from the start.
Our proof is based on the method of Lagrangian multipliers. Here we give only an outline and refer the interested reader to the Supplemental Material for the complete argument. First, write the program in terms of the Choi operator $C_{\mathcal{R}}$ of $\mathcal{R}$. Recalling that the reverse process goes from $B$ back to $A$, the trace-preserving constraint becomes $\mathrm{Tr}_{A}[C_{\mathcal{R}}]-\mathds{1}_{B}=0$. Then, define the Lagrangian of the problem as
@@
$C_{\mathcal{R}}=\left(D^{*}\otimes\sqrt{\gamma}\right)C_{\mathcal{E}}^{T}\left(D^{T}\otimes\sqrt{\gamma}\right)\,,$ (16)
where $D^{*}$ is the complex conjugate of the operator $D$ defined in Eq. 13, computed, like the transpose, with respect to the basis fixed by the Choi operator. In the Supplemental Material, we show that Eq. 16 and the value of $\Lambda$ given by

### v3

# Quantum Bayes’ rule and Petz transpose map from the minimum change principle

@@ Abstract
Bayes’ rule, which is routinely used to update beliefs based on new evidence, can be derived from a principle of minimum change. This principle states that updated beliefs must be consistent with new data, while deviating minimally from the prior belief. Here, we introduce a quantum analog of the minimum change principle and use it to derive a quantum Bayes’ rule by minimizing the change between two quantum input-output processes, not just their marginals. This is analogous to the classical case, where Bayes’ rule is obtained by minimizing several distances between the joint input-output distributions. When the change maximizes the fidelity, the quantum minimum change principle has a unique solution, and the resulting quantum Bayes’ rule recovers the Petz transpose map in many cases.
@@
Introduction.—Usually demonstrated by simple counting arguments involving urns and balls, Bayes’ rule has actually been argued to play a much deeper role in probability theory and logic, as the only consistent system for updating one’s beliefs in light of new evidence [1, 2, 3, 4, 5, 6]. As an alternative to the above axiomatic approach, Bayes’ rule can also be derived from a variational argument: the updated belief should be consistent with the new observations while deviating as little as possible from the initial belief. This is known as the minimum change principle [7, 8, 9, 10]. It formalizes the intuition that the new information should be incorporated into the agent’s knowledge in the “least committal” way, e.g. without introducing biases unwarranted by the data. Such fundamental insights can be seen as at least a motivation, if not an explanation, for the extraordinary effectiveness of Bayesian statistical inference in virtually all areas of knowledge.
@@
If one considers quantum theory as a noncommutative extension of probability theory, one would expect that there would also be a sound analog of Bayes’ rule. However, the status of a quantum Bayes’ rule is still much debated, with many alternatives, often inequivalent to each other, having been proposed in the past decades [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Among these proposals, the Petz transpose map [26, 27] stands out as the only quantum Bayes’ rule that satisfies a set of axioms analogous to the classical Bayes’ rule [25].
@@
Attempts have also been made to derive a quantum Bayes’ rule from optimizations involving the “posterior” state. For instance, Ref. [19] minimizes a loss function related to the quantum relative entropy, while Refs. [28, 22] optimize an upper bound on a distance measure between two estimators of the initial and final states. However, these methods, while involving optimization, focus primarily on the marginals of the process rather than on the entire process. As a result, they do not fully reflect the minimality of the change over the entire process, which is instead the core argument from which Bayes’ rule and its generalizations, such as Jeffrey’s theory of probability kinematics, emerge.
@@
In this work, we take a decisive step towards solving this problem by presenting a new approach to the quantum Bayes’ rule, based on a natural quantum analog of the minimum change principle, which involves the entire process, not just its marginals (see Fig. 1 for a schematic representation). Specifically, when the change maximizes the quantum fidelity [29, 30], the resulting quantum Bayes’ rule can be derived analytically and corresponds to the Petz transpose map in many cases. This connection further strengthens the link between Bayes’ rule, the minimum change principle, and the Petz transpose map, thus justifying their broader applications in quantum information theory and possibly beyond.
@@
Classical Bayes’ rule from the minimum change principle.—Consider two random variables $X$ and $Y$ taking values in two finite alphabets, $\mathcal{X}=\{x\}$ and $\mathcal{Y}=\{y\}$, respectively. Assume that the initial belief about $X$ is represented by a distribution $\gamma(x)$, while the correlation between $X$ and $Y$ is given by a conditional probability distribution $\varphi(y|x)$. We can think of the latter as the “forward process” from $X$ to $Y$. Now, suppose we observe a certain value $Y=y_{0}$ and want to update our belief about $X$ in light of this new information. Bayes’ rule provides the formula for the posterior probability as:
@@
But how should the above be generalized to the situation where the new information does not come in the form of a definite value $y_{0}$ for $Y$, but as “soft evidence,” i.e., a probability distribution $\tau(y)$? As Jeffrey [31] argues, this is in fact the rule rather than the exception, since it is very rare that we can ascribe probability one to only one event $Y=y_{0}$, ruling out all others with absolute certainty. Jeffrey’s probability kinematics [31] and Pearl’s virtual evidence method [32] agree [33] that the correct way to proceed is to promote Bayes’ posterior to a full-fledged “reverse process” $\hat{\varphi}_{\gamma}(x|y)$ that yields the probabilities of $X$ given $Y$, so that the updated belief given the new evidence $\tau(y)$ becomes
@@
While Jeffrey’s and Pearl’s derivations of Eq. 2 rely on certain natural assumptions about the logic model or underlying Bayesian network, an alternative way to obtain Eq. 2 is through the minimum change principle [8, 7, 9, 10]. To use this principle, we consider the joint distribution of $X$ and $Y$ in the forward process, denoted as $P_{\rm fwd}(x,y):=\varphi(y|x)\gamma(x)$, and try to update it based on a subsequent observation of $Y$. The comparison will be done not only on the marginal distribution on $X$, but on the whole joint distribution $P_{\rm fwd}(x,y)$. The minimum change principle then determines the updated belief about the joint distribution, denoted as $P_{\rm rev}(x,y)$.
Concretely, the minimum change principle can be described as the minimization of a divergence measure between the prior belief $P_{\rm fwd}$ and the updated belief $P_{\rm rev}$ under the constraint that $P_{\rm rev}$ is compatible with the new observation $\tau(y)$. This can be expressed as the following optimization problem:
@@
where $\mathbb{D}\left(P_{\rm fwd},P_{\rm rev}\right)$ is a statistical divergence measure between the prior and posterior joint distributions. Common choices include the Kullback–Leiber divergence [7, 10], other $f$-divergences including Pearson divergence and Hellinger distance [34], zero-one loss [35], or the mean-square error of an estimation [36, 37]. In all these cases, the optimal value is achieved by
@@
The above equation shows that the minimum change principle applied to the joint distributions of $X$ and $Y$ leads to the Bayes–Jeffrey rule in Eqs. 1 and 2.
@@
Stochastic processes in the quantum case correspond to completely positive trace-preserving (CPTP) linear maps, also known as quantum channels. Following the classical Bayes’ rule, we start with a forward quantum process $\mathcal{E}$, i.e. a channel from system $A$ to system $B$. The goal is to determine a backward quantum process $\mathcal{R}$, which is again a channel, but going from $B$ to $A$. Note that the adjoint map $\mathcal{E}^{\dagger}$, defined by trace duality as $\operatorname{Tr}\!\left[\mathcal{E}(X)\;Y\right]=:\operatorname{Tr}\!\left[X\;\mathcal{E}^{\dagger}(Y)\right]$ for all operators $X$ and $Y$, is a completely positive linear map from $B$ to $A$, although it is generally not trace-preserving and thus not a channel. As in the classical case, we assume a prior belief about the input, but instead of a probability distribution, the prior belief is now given as a quantum state $\gamma\in\mathcal{S}({\mathcal{H}}_{A})$  [^1] There are also proposals of quantum Bayes’ rule where the belief is described by objects other than a density matrix, such as a distribution over states [71], or a joint state over the system of interest and a reference system [72]. We leave their relation to the minimum change principle open for future studies..
Next, we look for a quantum analog of the joint input-output distribution of a stochastic process, such as the classical $P_{\rm fwd}(x,y)=\varphi(y|x)\gamma(x)$, which explicitly encodes correlations between inputs and outputs. In the classical case, the joint input-output distribution is defined directly by conditional probabilities, whereas in the quantum case, channels map input states to output states without providing a direct analog of a joint distribution.
To overcome this limitation and to represent the input-output correlations of a quantum channel, we adopt a purification-based construction. Specifically, we consider the canonical purification of $\gamma$ [39, 40]
[…]
where $A_{1}$ and $A_{2}$ are two copies of system $A$ and {|$\{\ket{i}\}$ is a chosen orthonormal basis of ${\mathcal{H}}_{A}$. This pure state has the following marginal states
$\mathrm{Tr}_{A_{2}}[|\sqrt{\gamma}\rangle\!\rangle\langle\!\langle\sqrt{\gamma}|]=\gamma\,,\qquad\mathrm{Tr}_{A_{1}}[|\sqrt{\gamma}\rangle\!\rangle\langle\!\langle\sqrt{\gamma}|]=\gamma^{T}\,,$ (7)
where the superscript notation $\bullet^{T}$ denotes the transposition taken with respect to the orthonormal basis {|$\{\ket{i}\}$.
Now, since we have two (correlated) copies of the input system, one could apply $\mathcal{E}$ to the first system and obtain $\mathcal{E}(\gamma)$, while the second system remains in state $\gamma^{T}$. Explicitly, this results in the state
$\begin{split}Q_{\rm fwd}&:=(\mathcal{E}\otimes\mathcal{I})\left(|\sqrt{\gamma}\rangle\!\rangle\langle\!\langle\sqrt{\gamma}|\right)\\ &=\left(\mathds{1}_{B}\otimes\sqrt{\gamma^{T}}\right)C_{\mathcal{E}}\left(\mathds{1}_{B}\otimes\sqrt{\gamma^{T}}\right),\end{split}$ (8)
where $\mathcal{I}$ is the identity map on system $A_{2}$, $\mathds{1}_{B}$ is the identity operator on system $B$ and $C_{\mathcal{E}}=\sum_{i,j}\mathcal{E}(|i\rangle\!\!\!\;\langle j|)\otimes|i\rangle\!\!\!\;\langle j|\in\mathcal{L}_{+}({\mathcal{H}}_{B}\otimes{\mathcal{H}}_{A})$ is the Choi operator of the forward process $\mathcal{E}$ [41]. Computing the marginal operators we obtain
@@
Such a representation of a quantum process has appeared in the literature in a context comparing quantum processes to classical conditional probabilities [42, 43]. A benefit of this representation is that $Q_{\rm fwd}$ is always in $\mathcal{S}({\mathcal{H}}_{B}\otimes{\mathcal{H}}_{A})$, i.e., it is a bipartite state for all choices of channel $\mathcal{E}$ and prior state $\gamma$, and many divergence functions are well-defined and possess nice properties when evaluated on states.
Although the operator $Q_{\rm fwd}$ is not a quantum state over time (the marginal on the input is not the input but its transpose; see Refs. [44, 24] for details), it is nonetheless very close in spirit to a state over time. In the fully commuting case, it reduces to the conventional joint input-output probability distribution $P_{\rm fwd}(x,y)$. For this reason, we will borrow the notation from Ref. [20] and define $\mathcal{E}\star\rho$ as
$\mathcal{E}\star\rho:=\left(\mathds{1}_{B}\otimes\sqrt{\rho^{T}}\right)C_{\mathcal{E}}\left(\mathds{1}_{B}\otimes\sqrt{\rho^{T}}\right),$ (10)
so that our forward process operator becomes
$Q_{\rm fwd}=\mathcal{E}\star\gamma\;.$ (11)
@@
$Q_{\rm rev}:=(\mathcal{R}\star\tau)^{T}=(\sqrt{\tau}\otimes\mathds{1}_{A})\;C_{\mathcal{R}}^{T}\;(\sqrt{\tau}\otimes\mathds{1}_{A})\;,$ (12)
@@
The quantum minimum change principle.—We are now ready to formulate the quantum minimum change principle in a rigorous way. Here, we choose to measure “change” in terms of the (square-root) fidelity, defined as [29, 30]
@@
The quantum minimum change principle, similar to Eq. 5, minimizes the deviation between $Q_{\rm fwd}=\mathcal{E}\star\gamma$ and $Q_{\rm rev}=(\mathcal{R}\star\tau)^{T}$, or equivalently, maximizes their fidelity. The principle is formally expressed as follows:
$\begin{split}\max_{\mathcal{R}}\quad&F\left(\mathcal{E}\star\gamma,(\mathcal{R}\star\tau)^{T}\right)\\ \text{subject to}\quad&\mathcal{R}\text{\rm~is CPTP}\;.\end{split}$ (15)
@@
An example comparing the Petz transpose map and the optimal solution of Eq. 15 is shown in Fig. 2.
Figure 2: Example comparing the Petz transpose map with the optimal map obtained from the minimum change principle. The forward map $\mathcal{E}$ models one step of a thermalizing quantum machine where the system interacts with a thermal state $\xi$ [46]. Concretely, we consider qubit systems, and the map is defined as $\mathcal{E}(\rho):=\mathrm{Tr}_{A^{\prime}}[V(\rho\otimes\xi_{A^{\prime}})V^{\dagger}]$, where $V:=\cos\frac{\pi}{8}~\mathds{1}+i\sin\frac{\pi}{8}~U_{\rm sw}$ is the partial swap operator with angle $\frac{\pi}{8}$, $U_{\rm sw}$ is the SWAP gate, and $\xi=0.95|0\rangle\!\!\!\;\langle 0|+0.05|1\rangle\!\!\!\;\langle 1|$. The prior belief is parameterized as $\gamma=(1-p)|0\rangle\!\!\!\;\langle 0|+p|1\rangle\!\!\!\;\langle 1|$, and the reference output state $\tau$ is chosen as $\tau=H\xi H$, where $H$ is the Hadamard gate. The figure plots the value of the objective function of Eq. 15 over $p\in[0.001,0.999]$, for $\mathcal{R}$ chosen to be the Petz transpose map [Eq. 18, orange dashed line] or the optimal map [Eq. 16, blue solid line]. In most cases $[\tau,\mathcal{E}(\gamma)]\neq 0$, and the Petz transpose map is suboptimal; while at $p=1.85-0.9\sqrt{2}\approx 0.577$, where $\mathcal{E}(\gamma)=\mathds{1}/2$ and $[\tau,\mathcal{E}(\gamma)]=0$, the Petz map achieves optimality.
Before fleshing out the proof of the above theorem (the full details can be found in Supplemental Material  [^2] See Supplemental Material for proofs the main theorem and lemmas, which includes Refs. [73, 74, 75, 76, 77]), let us first make some comments. First, it is easy to verify that the map given in Eq. 16 is indeed completely positive and trace-preserving. Complete positivity holds because $\mathcal{R}$ is the composition of three completely positive maps, that is, $D\bullet D^{\dagger}$, $\mathcal{E}^{\dagger}(\bullet)$, and $\sqrt{\gamma}\bullet\sqrt{\gamma}$. The trace-preservation condition is $\mathrm{Tr}[\mathcal{R}(\sigma)]=\mathrm{Tr}[D^{\dagger}\mathcal{E}(\gamma)D\sigma]=\mathrm{Tr}[\sigma]$ for any $\sigma$, equivalent to the condition
@@
    when $\mathcal{E}$ is a quantum-classical channel, e.g., a measurement channel, and the new information represented by $\tau$ is classical information about the measurement results.
The last condition makes Theorem 1 especially compelling in the context of observational entropy and the corresponding second law [50, 45, 51].
Third, when the commutativity condition $[\tau,\mathcal{E}(\gamma)]=0$ is satisfied and the retrodictive channel does not depend on $\tau$, our solution is in line with Jeffrey’s framework of probability kinematics [31], where the update rule relies solely on the prior $\gamma$ and the forward map $\varphi$, independently of the new information $\tau$. However, when $\mathcal{E}(\gamma)$ and $\tau$ do not commute, the quantum case departs from Jeffrey’s framework, and the role of $\tau$ becomes interpretationally significant. If $\tau$ is taken as a subjective prior belief about what the observer expected to observe before obtaining the actual outcome, then its sole purpose is to “guide” the optimization according to the principle of minimum change. In this case, the update rule is given by Eq. 16 with $\sigma$ representing the newly acquired information, and the rule remains independent of $\tau$. In contrast, if $\tau$ is identified with the actual observed evidence—namely, if one sets $\sigma=\tau$ in Eq. 16—then the quantum counterpart of Jeffrey’s retrodicted distribution $\gamma^{\prime}(x)$ from Eq. 2 becomes the quantum state
@@
While the CPTP map $\mathcal{R}$ is always linear, the overall dependence of $\gamma^{\prime}$ on $\tau$ is not, in contrast to the classical case. Again, it is the commutativity condition $[\tau,\mathcal{E}(\gamma)]=0$ that restores the linear dependence of $\gamma^{\prime}$ on $\tau$, and thus in the classical case the two interpretations make no difference. The functional dependence of the retrodicted state on the new information is an important aspect of the problem, which we leave open for future research, especially in view of the axiomatic approach proposed in [25], where linearity of the retrodiction is assumed from the start.
Fourth, when $[\tau,\mathcal{E}(\gamma)]=0$, Theorem 1 can also deal with a non-invertible $\tau$: in this case, we can replace $\tau$ with $\tau_{\epsilon}:=(1-\epsilon)\tau+\epsilon u$, where $\epsilon>0$ and $u$ denotes the uniform (i.e. maximally mixed) state; find the optimal reverse channel, which does not depend on $\tau_{\epsilon}$ as the latter preserves the commutation relation; and finally take the limit for $\epsilon\to 0$. In this way, we see that Theorem 1 does indeed contain the conventional Bayes’ rule (1), where the new information is given as a delta function.
Our proof is based on the method of Lagrangian multipliers. Here, we give only an outline and refer the interested reader to Appendix A for the complete argument. First, write the program in terms of the Choi operator $C_{\mathcal{R}}$ of $\mathcal{R}$. Recalling that the reverse process goes from $B$ back to $A$, the trace-preserving constraint becomes $\mathrm{Tr}_{A}[C_{\mathcal{R}}]-\mathds{1}_{B}=0$. Then, define the Lagrangian of the problem as
@@
$C_{\mathcal{R}}=\left(D^{T}\otimes\sqrt{\gamma}\right)C_{\mathcal{E}}^{T}\left(D^{*}\otimes\sqrt{\gamma}\right)\,,$ (20)
where $D^{*}$ is the complex conjugate of the operator $D$ defined in Eq. 17, computed with respect to the same basis as the transpose. In Appendix A, we show that Eq. 20 and the value of $\Lambda$ given by